• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name R. Michaels

    Log entry time 15:45:34 on October16,2009

    Entry number 295239


    keyword=CH scaler problems (again)

    Investigated the scaler in CH crate (ROC23).
    Note, this is involved in CH synch check and in Hall C BCM.
    The value does not change.  The control lights don't
    blink, meaning it's not accepting LNE pulses.  But LNE
    pulses and gate (GMN2) are there at the inputs.  We
    may try reseting the crate or restarting DAQ.  This was
    a problem earlier in the experiment too.  We thought we
    fixed it and -- for a long time -- it was not observed.